OpenStudy (anonymous):

Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. (Hint: A chord divides a circle into two segments. In problem 1, you found the area of the smaller segment.)

OpenStudy (anonymous):

the answer needs to be like this: \[\frac{ ??? }{ ? }\pi + ??\sqrt{?}\]